# Large Language Models in Multi-Round, Multi-Party Dialogues

### v1

@@ 3.4  DICE-Score
Alignment with Human Evaluation.
The proposed metric, DICE-Score, was developed to quantify task difficulty across a dataset of 1,607 samples and was validated through human evaluation using a statistically grounded subset of 311 samples. This subset size was determined based on a 95% confidence level, a 5% margin of error, and a conservative estimate of maximum variability ($p=0.5$). The calculation incorporated a Finite Population Correction (FPC) to account for the dataset’s finite size.
@@
To assess the alignment between human performance and the proposed difficulty metric, we computed the Pearson correlation coefficient. The analysis revealed a strong negative correlation ($r\approx-0.984$), indicating that higher DICE-Score values were associated with lower human accuracy. This trend is consistent across rounds, from 80.5% accuracy at $DICE=1.42$ (Round 1) to 49.3% at $DICE=5.36$ (Round 4). A t-test confirmed the statistical significance of this correlation, yielding a t-value of approximately $-8.38$ ($p<0.01$, 2 degrees of freedom).
@@
Round $N$ Acc (%) DICE-Score
@@
Table 4: Human Evaluation Results by Round. Accuracy denotes the proportion of correctly answered DICE-Bench  samples by human participants. $N$ refers to the sample size per round.
@@ 4.1 Model Selection
We evaluated a total of 19 LLMs that support at least 8k context window size in DICE-Bench, both closed-source and open-source LLMs. The closed-source cohort includes GPT-4o and GPT-4o-mini OpenAI et al. (2024), along with Gemini 2 Flash and Gemini 2 Flash Lite Team et al. (2020). Meanwhile, our open-source lineup spans a wide range of general-purpose models, including LLaMA3 Touvron et al. (2023), Qwen2.5 Qwen et al. (2025), Mistral Jiang et al. (2023), EXAONE Research et al. (2024), Phi4 Abdin et al. (2024), GLM4-Chat GLM et al. (2024). In addition, we evaluate tool-specific models that have been fine-tuned on tool datasets, including Hammer2.1 Wang et al. (2024a), ToolAce Liu et al. (2024), CALM Acikgoz et al. (2025), NexusRaven-V2 team (2023), Granite Abdelaziz et al. (2024).
@@ 4.3.2 Analysis
 DICE-Score Validity.
DICE-Score is designed to quantify how dispersed the critical information is in multi-round dialogues, thereby indicating the difficulty of function-calling tasks. Our experiments provide strong evidence for its validity. As demonstrated in Table 5, model performance steadily declines as the number of rounds increases, suggesting that the task becomes more challenging when essential information is spread out. In parallel, Table 3 shows that DICE-Score rises with each additional round. In DICE-Score, a higher value indicates that the crucial details are more sparsely distributed across the dialogue, directly correlating with the increased difficulty of retrieving that information. This inverse relationship, where an increase in DICE-Score corresponds with a drop in performance, supports the effectiveness of our metric in capturing task complexity. In essence, the consistent alignment between higher DICE-Score values and reduced model performance confirms the validity of DICE-Score  as a reliable measure of the challenges inherent in function-calling tasks.
Uncovering True Performance Factors
Figure 7 reveals that as the number of rounds increases, model performance declines sharply, especially when moving from Round 1 to Round 2. While this decrease might be partially attributed to the accumulation of dialogue, introducing long-context challenges, other factors could also be contributing.
To dissect the causes of increased task difficulty in longer dialogues, we employ the DICE-Score. By incorporating a logarithmic transformation in its numerator, the DICE-Score prevents task difficulty from being overly influenced by merely longer utterances, allowing us to isolate other factors.
As shown in Table 3, the DICE score consistently increases with each additional round, indicating that the challenge is not simply due to the long-context problem in LLMs, but rather stems from the fact that the essential information for function-calling becomes limited and sporadically distributed as the dialogue lengthens. This suggests that the primary difficulty in function-calling lies in retrieving crucial, dispersed information from dialogues with multiple utterances.
@@ B.2 Correlation Analysis
$t\approx\frac{-0.984\cdot\sqrt{2}}{\sqrt{1-0.968}}=\frac{-1.391}{0.166}\approx-8.38$ (9)
With 2 degrees of freedom, this result is statistically significant ($p<0.01$), confirming a strong negative correlation between human accuracy and task difficulty as measured by DICE-Score.

### v2

@@ 3.4  DICE-Score
Alignment between DICE-Score and Human Performance.
To validate DICE-Score, we measured the Spearman Correlation between the human performance on DICE-Bench and DICE-Score using a statistically grounded subset of 311 samples. This subset size was determined based on a 95% confidence level, a 5% margin of error, and a conservative estimate of maximum variability ($p=0.5$). The calculation incorporated a Finite Population Correction (FPC) to account for the dataset’s finite size.
@@
To assess the alignment between human performance and the proposed difficulty metric, we computed the Pearson correlation coefficient. The analysis revealed a strong negative correlation ($r\approx-0.984$), indicating that higher DICE-Score values were associated with lower human accuracy. This trend is consistent across rounds, from 80.5% accuracy at $DICE=1.42$ (Round 1) to 49.3% at $DICE=5.36$ (Round 4). A t-test confirmed the statistical significance of this correlation, yielding a t-value of approximately $-7.81$ ($p<0.05$, 2 degrees of freedom).
@@
Round $N$ EM (%) DICE-Score
@@
Table 4: Human Evaluation Results by Round. EM (%) denotes the Human performance using EM metrics on DICE-Bench  samples. $N$ refers to the sample size per round. DICE-Scorecolumn refers to the DICE-Score for each round on the entire dataset.
@@ 4.1 Model Selection
We evaluated a total of 19 LLMs that support at least 8k context window size. Also we excluded the reasoning models. The closed-source cohort includes GPT-4o and GPT-4o-mini OpenAI et al. (2024), along with Gemini 2 Flash and Gemini 2 Flash Lite Team et al. (2020). Meanwhile, our open-source lineup spans a wide range of general-purpose models, including LLaMA3 Touvron et al. (2023), Qwen2.5 Qwen et al. (2025), Mistral Jiang et al. (2023), EXAONE Research et al. (2024), Phi4 Abdin et al. (2024), GLM4-Chat GLM et al. (2024). In addition, we evaluate tool-specific models that have been fine-tuned on tool datasets, including Hammer2.1 Wang et al. (2024a), ToolAce Liu et al. (2024), CALM Acikgoz et al. (2025), NexusRaven-V2 team (2023), Granite Abdelaziz et al. (2024).
@@ 4.3.1 Results
Table 5 also shows that tool-specific models such as ToolAce-8B, CALM-8B, NexusRaven-V2-13B, and Granite-20B show poor performance compared to the other models. We assume that since those tool-specific models are finetuned to accept a single instruction, it is not generalized to accept the multi-party dialogues as the input.
@@ 4.3.2 Analysis
Validity and Performance Analysis.
Our analysis reveals that the observed performance decline in multi-round dialogues is not primarily due to increased input length (long-context limitations), but rather due to the dispersion of critical tool-related information across dialogue rounds, a factor effectively captured by DICE-Score. Evidence from Table 5 and Figure 4 demonstrates a clear inverse correlation between DICE-Score values and model performance. Specifically, because the numerator of DICE-Score employs a logarithmic scale, it effectively isolates information dispersion from utterance length, confirming that retrieving sparse and fragmented tool-related details significantly impacts model outcomes. This consistent relationship across datasets and models substantiates the validity of DICE-Score as a measure of information dispersion and task complexity.
Influence of Dialogue Type.
Our analysis shows that function-calling performance varies significantly depending on dialogue type, with the $Eristic$ type notably exhibiting lower performance compared to others due to frequent stance changes among speakers. As illustrated in the middle lower plot of Figure 7, dialogue types such as $Persuasion\ Deliberation\&Negotiation$, and $Inquiry\&Information\ Seeking$ follow similar performance trends, whereas $Eristic$ dialogues consistently yield lower EM scores. Qualitative analysis attributes this disparity to the inherent complexity of $Eristic$ dialogues, where frequent shifts in speaker positions and their opinion create ambiguity, complicating the task for LLMs when identifying relevant information for accurate function calling.
Effect of Fine-tuning Objectives on Model Performance.
Our analysis demonstrates that the performance of LLMs on DICE-Bench is significantly influenced by their fine-tuning objectives, with general conversational models outperforming models specifically fine-tuned for single-turn function calling. For example, function-calling-specific models like ToolAce-8B, CALM-8B, and NexusRaven-V2-13B exhibit notably poorer performance compared to general-purpose models. In contrast, the GLM4-9B-Chat model, optimized for multi-turn dialogue understanding, achieves superior performance even relative to larger models. This disparity arises because single-turn function-calling datasets, used to fine-tune specialized models, inadequately represent the complexities of realistic multi-round interactions encountered in DICE-Bench. Consequently, models trained on conversational tasks, such as GLM4-9B-Chat, are better suited for handling these more complex, multi-turn function-calling scenarios.
@@ B.2 Correlation Analysis
$t\approx\frac{-0.984\cdot\sqrt{2}}{\sqrt{1-0.968}}=\frac{-1.391}{0.179}\approx-7.81$ (9)
With 2 degrees of freedom, this result is statistically significant ($p<0.05$), confirming a strong negative correlation between human accuracy and task difficulty as measured by DICE-Score.